# Elephant random walks on infinite Cayley trees

### v2

@@ Abstract.
Elephant random walk; Bethe lattice; Cayley trees; Pólya’s urn
@@ 1. Introduction
[BL19] and [Ber22] studied the extension of the ERW to $\mathbb{Z}^{d}$, also called the multi-dimensional elephant random walk (MERW), and established various limit theorems. Recent studies, such as [Qin25] and [CL24], have investigated recurrence and transience in MERW. [Qin25] established that for $d\geq 3$, the walk is transient, whereas for $d=2$, the walk becomes transient for $p\geq 5/8$. Recurrence in the sub-critical regime $p<5/8$ for $d=2$ was also established by [CL24] by a rough comparison with the standard plane random walk.
@@
Beyond the classical lattice setting, random walks have been extensively studied on a variety of infinite graphs. Indeed, a fruitful direction of modern research in probability has been the investigation of how the large-scale geometric properties of a space are reflected in the behaviour of random walks on that space. Regular infinite graphs arise naturally as Cayley graphs of finitely generated groups. An important example is the infinite $d$-ary tree $\mathbb{T}_{d}$ (also called the Bethe lattice in the physics literature), which arises as the Cayley graph, with respect to the standard generators, of $\mathbb{Z}^{*d_{1}}*\mathbb{Z}_{2}^{*d_{2}}$, with $d=2d_{1}+d_{2}$, where * denotes the free product of groups. The simple (i.e. nearest neighbour) random walk (SRW) on $\mathbb{T}_{d}$ is known to be ballistic for $d\geq 3$, with speed $\frac{d-2}{d}$, and, a fortiori, transient (see, e.g., Chapter 13 of [LP17]). In general, the study of random walks on groups has uncovered profound connections between probability theory and geometric group theory.
@@
Motivated by this, in this article, we introduce elephant random walks on finitely generated groups, extending the notion of random walks with long-range memory beyond classical lattices, the goal being to explore how such memory interacts with the geometry of the group. As a first step in this broader program, we focus on the infinite $d$-ary tree $\mathbb{T}_{d}$ for $d\geq 3$, viewed as the Cayley graph of a finitely generated group, and lay the foundation for future investigations on more general groups and graphs. We show that memory has no impact on the asymptotic speed (and hence recurrence/transience) of the walk, thus matching the escape rate $\frac{d-2}{d}$ of the SRW. We also establish upper bounds on the rate of convergence to the asymptotic speed. These upper bounds exhibit a phase transition at $p=\frac{d+1}{2d}$, below which they are of order $n^{-1/2}$ and slower otherwise. Numerical experiments suggest that these upper bounds are tight. We also obtain estimates on the return probability and present several open problems.
@@
We would like to emphasize here that the non-commutative and non-Markovian nature of the ERW on a non-abelian group makes its analysis challenging. Indeed, due to the lack of the Markov property, a direct potential-theoretic approach is infeasible. For the same reason, it is not clear how tools from sub-additive ergodic theory could be applied. While our proofs are martingale-based, due to non-commutativity, they are of a rather different flavour from the (also martingale-based) techniques used to analyse elephant random walks on $\mathbb{Z}^{d}$. Indeed, for the MERW, the position of the walker and hence their distance from the origin can be written succinctly in terms of the step counts along the various axis directions in the past (see, e.g., Eq. (13) of [Qin25]). Such a simple representation is, however, not possible on non-abelian groups, where the geometry of the Cayley graph enters non-trivially, interacting with step counts in a more intricate manner. It is necessary to decouple the urn process governing the step counts from the geometry of the Cayley graph in order to make progress.
@@
The rest of the paper is organised as follows. In Section 2, we describe the model and state our main results. In Section 3, we set up some preliminaries and prove a crucial concentration estimate. Finally, Section 4 collects the proofs of our main results.
@@ 2.1. Random walk on finitely generated groups
We first briefly recall the definition of random walks on groups.
Let $\mu$ be a probability measure on a group $\Gamma$. The (left-) $\mu$-random walk on $\Gamma$ is the (Markovian) stochastic process $(g_{n}\cdots g_{1})_{n\geq 1}$, where the “increments” $g_{i}$ are i.i.d. random elements distributed as $\mu$.
@@ 2.2. The ERW on finitely generated groups
The ERW is a special case of the so-called step-reinforced random walks (SRRW). [Kim14] studied an SSRW on $\mathbb{Z}$ which is equivalent to the ERW (this was pointed out by [Kür15]; see also [Kür16]). Since then SSRWs with more general step distributions have been studied by various authors, including [Bus18, Ber20b, Ber20a, Ber21b, Ber21a, BRO22, HZ24, Qin25]. One may readily generalise the positively step reinforced random walk (pSSRW) to groups as follows: Given a measure $\mu$ on a group $\Gamma$, we let $g_{1}\sim\mu$. Given steps $g_{1},\ldots,g_{n}$, at time $n+1$, we
@@ 2.3. ERW on Cayley trees
As stated in the introduction, the goal of this paper is to study ERW on finitely generated non-abelian infinite groups, by studying perhaps the simplest case where the Cayley graph is the infinite $d$-regular tree $\mathbb{T}_{d}$, with $d\geq 3$. Writing $\mathbb{Z}\cong\langle a,a^{-1}\rangle$ and $\mathbb{Z}_{2}\cong\langle b\mid b^{2}=e\rangle$, one may realise $\mathbb{T}_{d}$ as a Cayley graph by considering the free product $\mathbb{Z}^{*d_{1}}*\mathbb{Z}_{2}^{*d_{2}}$, with $d=2d_{1}+d_{2}$. Indeed, $\Gamma=\mathbb{Z}^{*d_{1}}*\mathbb{Z}_{2}^{*d_{2}}$ is generated by $\mathcal{S}=\{a_{1},a_{1}^{-1},\ldots,a_{d_{1}},a_{d_{1}}^{-1}\}\cup\{b_{1},\ldots,b_{d_{2}}\}$, where the only non-trivial relations satisfied by these generators are $b_{j}^{2}=e$ for $j=1,\ldots,d_{2}$.
@@
We shall use the notations $A_{n}\lesssim B_{n}$ or $B_{n}\gtrsim A_{n}$ to mean that there is some constant $C>0$, not depending on $n$, such that $A_{n}\leq CB_{n}$ for all $n\geq 1$. The constant $C$ will typically depend on various parameters such as $d,p,m$, etc., which will be clear from the context. The notation $A_{n}\asymp B_{n}$ shall mean that both $A_{n}\lesssim B_{n}$ and $B_{n}\lesssim A_{n}$ hold true. For random variables, we use the standard $O_{P}$ and $o_{P}$ notations.
@@ 2.5. Main results
For $m=2$, our methods give an extra (likely suboptimal) $\log n$ factor in the subcritical and critical regimes:
[…]
@@
Remark 2.3.
For $m>2$, our proof technique only gets us the following:
[…]
@@
In the course of proving Theorem 2.2, we also obtain estimates on the probability of returning to $e$ at time $n$. This is known to be exponentially decaying in $n$ for the SRW (i.e. $p=1/d$). Our bounds decay exponentially in $n$ for $0\leq p<1/2$ (except for the special case $(p,d)=(0,3)$). For $p\geq 1/2$, or $(p,d)=(0,3)$, we only get linear decay in $n$.
@@
Let $p\in[1/2,1)$ or $(p,d)=(0,3)$. Then
$\mathbb{P}(\Delta_{n}=0)\lesssim n^{-1}.$
Remark 2.4.
The linear decay in the latter regime is suboptimal as the following example shows. For $d\geq 2$ and $\Gamma=\mathbb{Z}^{*d}$, the free group of rank $d$, we have $\|S_{n}\|_{1}\leq\Delta_{n}$ , where $S_{n}$ is the MERW on $\mathbb{Z}^{d}$ and $\|\cdot\|_{k}$, for $k\geq 1$, denotes the $\ell^{k}$ norm. This is due to the fact that if $a$ is in the generating set for $\Gamma$, then $a$ and $a^{-1}$ cancel in a sample path only when they appear next to each other. As such, from the proof of Proposition 3.5 of [Qin25], we get that
$\mathbb{P}(\Delta_{n}=0)\leq\mathbb{P}(\|S_{n}\|_{1}=0)=\mathbb{P}(\|S_{n}\|_{2}=0)\lesssim n^{-\frac{d}{2}}.$
@@
An inspection of the proof of Theorem 2.3 reveals that one would get (streched) exponential decay of the return probability $\mathbb{P}(\Delta_{n}=0)$ for $p\geq 1/2$, or for the special case $(p,d)=(0,3)$, if one is able to establish (stretched) exponential concentration inequalities for $\Xi_{n}$ in these regimes.
@@
Prove (stretched) exponential decay of the return probability in all regimes of $p$.
@@
For SRW, i.e. $p=1/d$, Theorem 2.4 reduces to the well-known central limit theorem: $\sqrt{n}(\frac{\Delta_{n}}{n}-\frac{d-2}{d})\xrightarrow{d}N(0,\frac{4(d-1)}{d^{2}})$.
Theorem 2.4 implies that in the critical and the supercritical regimes, one has
@@
Our simulations suggest that the limits are Gaussian in the subcritical and critical regimes. In fact, in the subcritical regime the limit appears to be $N(0,\frac{4(d-1)}{d^{2}})$, regardless of the value of $p$ (see Figure 3). In view of Theorem 2.4, this suggests that $\sqrt{n}\cdot\Xi_{n}=o_{p}(1)$ in the subcritical regime. Our simulations seem to corroborate this (see Figure 4).
@@ 3.1. The urn connection
Let $p\in[0,1)$. Then for any $m\geq 1$ and $a\in\mathcal{S}$,
$\mathbb{E}\bigg|\frac{N_{n}(a)}{n}-\frac{1}{d}\bigg|^{m}\lesssim r_{n}^{-m}.$
For the MERW on $\mathbb{Z}^{d}$, a result of this form was worked out in Lemma 4.3 of [Qin25] for the moments of $\big|\frac{N_{n}(\mathbf{e}_{i})+N_{n}(-\mathbf{e}_{i})}{n}-\frac{1}{d}\big|$. The proof utilises a stochastic approximation decomposition together with the Burkholder-Davis-Gundy inequality [Dav70]. Since that argument works here almost verbatim, we omit the proof.
@@
Let $p\in[0,1)$. Then for $m\in[1,2)$, we have
$\mathbb{E}|\Xi_{n}|^{m}\lesssim r_{n}^{-m}.$
For $m=2$, we have
$\mathbb{E}|\Xi_{n}|^{2}\lesssim\begin{cases}r_{n}^{-2}\log n&\text{if }p\leq p_{d},\\ r_{n}^{-2}&\text{if }p>p_{d}.\end{cases}$
For $p\leq p_{d}$ and $m>2$, we have
$\mathbb{E}|\Xi_{n}|^{m}\lesssim n^{-1}.$
Finally, for $p>p_{d}$,
$\mathbb{E}|\Xi_{n}|^{m}\lesssim\begin{cases}r_{n}^{-m}\log n&\text{if }2<m\leq\frac{d-1}{d(1-p)},\\ n^{-1}\log n&\text{if }m=\frac{d-1}{d(1-p)},\\ n^{-1}&\text{if }m>\frac{d-1}{d(1-p)}.\end{cases}$
@@
$\bigg|\frac{1}{n}\sum_{k=0}^{n-1}\bigg(\frac{\ell_{k}}{k}-\frac{\mathds{1}(\Delta_{k}>0)}{d}\bigg)\bigg|^{m}$ $\leq\bigg(\frac{1}{n}\sum_{k=0}^{n-1}\bigg|\frac{\ell_{k}}{k}-\frac{\mathds{1}(\Delta_{k}>0)}{d}\bigg|\bigg)^{m}$
$\leq\frac{1}{n}\sum_{k=0}^{n-1}\bigg|\frac{\ell_{k}}{k}-\frac{\mathds{1}(\Delta_{k}>0)}{d}\bigg|^{m}$
$\leq\frac{1}{n}\sum_{k=1}^{n-1}\sum_{a\in\mathcal{S}}\bigg|\frac{N_{k}(a)}{k}-\frac{1}{d}\bigg|^{m}$
$=\sum_{a\in\mathcal{S}}\frac{1}{n}\sum_{k=1}^{n-1}\bigg|\frac{N_{k}(a)}{k}-\frac{1}{d}\bigg|^{m}.$
Then, using Lemma 3.1, we get that
$\mathbb{E}|\Xi_{n}|^{m}\lesssim\sum_{a\in\mathcal{S}}\frac{1}{n}\sum_{k=1}^{n-1}r_{k}^{-m},$
from which the desired upper bounds follow by applying Lemma 3.2 to the right-hand side. ∎

### v3

@@ Abstract.
Elephant random walk; step-reinforced random walk; Bethe lattice; Cayley trees; Pólya’s urn
@@ 1. Introduction
[BL19] and [Ber22] studied the extension of the ERW to $\mathbb{Z}^{d}$, also called the multi-dimensional elephant random walk (MERW), and established various limit theorems. Recent studies, such as [Qin25a] and [CL24], have investigated recurrence and transience in MERW. [Qin25a] established that for $d\geq 3$, the walk is transient, whereas for $d=2$, the walk becomes transient for $p\geq 5/8$. Recurrence in the sub-critical regime $p<5/8$ for $d=2$ was also established by [CL24] by a rough comparison with the standard plane random walk.
@@
Beyond the classical lattice setting, random walks have been extensively studied on a variety of infinite graphs. Indeed, a fruitful direction of modern research in both probability theory and statistical physics has been the investigation of how the large-scale geometric properties of a space are reflected in the behaviour of random walks on that space. Regular infinite graphs arise naturally as Cayley graphs of finitely generated groups. An important example is the infinite $d$-ary tree $\mathbb{T}_{d}$ (also known as the Bethe lattice in statistical physics, where it has been widely used to simplify problems that are difficult to solve on regular lattices), which arises as the Cayley graph, with respect to the standard generators, of $\mathbb{Z}^{*d_{1}}*\mathbb{Z}_{2}^{*d_{2}}$, with $d=2d_{1}+d_{2}$, where * denotes the free product of groups. The simple (i.e. nearest neighbour) random walk (SRW) on $\mathbb{T}_{d}$ is known to be ballistic for $d\geq 3$, with speed $\frac{d-2}{d}$, and, a fortiori, transient (see, e.g., Chapter 13 of [LP17]). In general, the study of random walks on groups has uncovered profound connections between probability theory and geometric group theory.
@@
Motivated by this, in this article, we introduce elephant random walks on finitely generated groups, extending the notion of random walks with long-range memory beyond classical lattices, the goal being to explore how such memory interacts with the geometry of the group. As a first step in this broader program, we focus on the infinite $d$-ary tree $\mathbb{T}_{d}$ for $d\geq 3$, viewed as the Cayley graph of a finitely generated group. We show that memory has no impact on the asymptotic speed (and hence recurrence/transience) of the walk, thus matching the escape rate $\frac{d-2}{d}$ of the SRW. We also establish upper bounds on the rate of convergence to the asymptotic speed. These upper bounds exhibit a phase transition at $p=\frac{d+1}{2d}$, below which they are of order $n^{-1/2}$ and slower otherwise. Numerical experiments suggest that these upper bounds are tight. We also obtain estimates on the return probability, which decays exponentially for any $0\leq p<1/2$ such that $(p,d)\neq(0,3)$. Based on extensive numerical experiments, we present several open problems on fluctuations around the limiting speed and the decay of the return probability in all memory regimes.
@@
We would like to emphasize here that the non-commutative and non-Markovian nature of the ERW on a non-abelian group makes its analysis challenging. Indeed, due to the lack of the Markov property, a direct potential-theoretic approach is infeasible. For the same reason, it is not clear how tools from sub-additive ergodic theory could be applied. While our proofs are martingale-based, due to non-commutativity, they are of a rather different flavour from the (also martingale-based) techniques used to analyse elephant random walks on $\mathbb{Z}^{d}$. Indeed, for the MERW, the position of the walker and hence their distance from the origin can be written succinctly in terms of the step counts along the various axis directions in the past (see, e.g., Eq. (13) of [Qin25a]). Such a simple representation is, however, not possible on non-abelian groups, where the geometry of the Cayley graph enters non-trivially, interacting with step counts in a more intricate manner. It is necessary to decouple the urn process governing the step counts from the geometry of the Cayley graph in order to make progress.
@@
The rest of the paper is organised as follows. In Section 2, we describe the model and state our main results. In Section 3, we set up some preliminaries and prove a crucial concentration estimate. Finally, Section 4 collects the proofs of our main results. Section 5 contains some concluding remarks.
@@ 2.1. Random walk on finitely generated groups
We first briefly recall the definition of random walks on (discrete) groups.
Let $\mu$ be a probability measure on a discrete group $\Gamma$. The (left-) $\mu$-random walk on $\Gamma$ is the (Markovian) stochastic process $(g_{n}\cdots g_{1})_{n\geq 1}$, where the “increments” $g_{i}$ are i.i.d. random elements distributed as $\mu$.
@@ 2.2. The ERW on finitely generated groups
The ERW is a special case of the so-called step-reinforced random walks (SRRW). [Kim14] studied an SSRW on $\mathbb{Z}$ which is equivalent to the ERW (this was pointed out by [Kür15]; see also [Kür16]). Since then SSRWs with more general step distributions have been studied by various authors, including [Bus18, Ber20b, Ber20a, Ber21b, Ber21a, BRO22, HZ24, Qin25b]. One may readily generalise the positively step reinforced random walk (pSSRW) to groups as follows: Given a probability measure $\mu$ on a group $\Gamma$, we let $g_{1}\sim\mu$. Given steps $g_{1},\ldots,g_{n}$, at time $n+1$, we
@@ 2.3. ERW on Cayley trees
As stated in the introduction, the goal of this paper is to study ERW on finitely generated non-abelian infinite groups, by studying perhaps the simplest case where the Cayley graph is the Bethe lattice, i.e. the infinite $d$-regular tree $\mathbb{T}_{d}$, with $d\geq 3$. Writing $\mathbb{Z}\cong\langle a,a^{-1}\rangle$ and $\mathbb{Z}_{2}\cong\langle b\mid b^{2}=e\rangle$, one may realise $\mathbb{T}_{d}$ as a Cayley graph by considering the free product $\mathbb{Z}^{*d_{1}}*\mathbb{Z}_{2}^{*d_{2}}$, with $d=2d_{1}+d_{2}$. Indeed, $\Gamma=\mathbb{Z}^{*d_{1}}*\mathbb{Z}_{2}^{*d_{2}}$ is generated by $\mathcal{S}=\{a_{1},a_{1}^{-1},\ldots,a_{d_{1}},a_{d_{1}}^{-1}\}\cup\{b_{1},\ldots,b_{d_{2}}\}$, where the only non-trivial relations satisfied by these generators are $b_{j}^{2}=e$ for $j=1,\ldots,d_{2}$.
@@
We shall use the notations $A_{n}\lesssim B_{n}$ or $B_{n}\gtrsim A_{n}$ to mean that there is some constant $C>0$, not depending on $n$, such that $A_{n}\leq CB_{n}$ for all $n\geq n_{0}$, for some $n_{0}\geq 1$. The constant $C$ will typically depend on various parameters such as $d,p,m$, etc., which will be clear from the context. The notation $A_{n}\asymp B_{n}$ shall mean that both $A_{n}\lesssim B_{n}$ and $B_{n}\lesssim A_{n}$ hold true. For random variables, we use the standard $O_{P}$ and $o_{P}$ notations.
@@ 2.5. Main results
In the supercritical regime, (4) holds for a larger range of $m$, namely, for all $m\in[1,\frac{d-1}{d(1-p)}]$.
@@
In the course of proving Theorem 2.2, we also obtain estimates on the probability of returning to $e$ at time $n$. This is known to be exponentially decaying in $n$ for the SRW (i.e. $p=1/d$). Our bounds decay exponentially in $n$ for $0\leq p<1/2$ (except for the special case $(p,d)=(0,3)$). For $p\geq 1/2$, or $(p,d)=(0,3)$, we only get stretched exponential decay in $n$.
@@
Let $p\in[1/2,1)$ or $(p,d)=(0,3)$. Then there exists a constant $C_{p,d}>0$, such that
$\mathbb{P}(\Delta_{n}=0)\lesssim\exp\big(-C_{p,d}\cdot r_{n}\big).$
@@
An inspection of the proof of Theorem 2.3 reveals that one would get exponential decay of the return probability $\mathbb{P}(\Delta_{n}=0)$ for $p\geq 1/2$, or for the special case $(p,d)=(0,3)$, if one is able to establish exponential concentration inequalities for $\Xi_{n}$ in these regimes.
@@
Prove exponential decay of the return probability in all regimes of $p$.
@@
For the SRW, i.e. $p=1/d$, Proposition 2.1 reduces to the well-known central limit theorem: $\sqrt{n}(\frac{\Delta_{n}}{n}-\frac{d-2}{d})\xrightarrow{d}N(0,\frac{4(d-1)}{d^{2}})$.
Proposition 2.1 implies that in the critical and the supercritical regimes, one has
@@
Our simulations suggest that the limits are Gaussian in the subcritical and critical regimes. In fact, in the subcritical regime the limit appears to be $N(0,\frac{4(d-1)}{d^{2}})$, regardless of the value of $p$ (see Figure 3). In view of Proposition 2.1, this suggests that $\sqrt{n}\cdot\Xi_{n}=o_{p}(1)$ in the subcritical regime. Our simulations seem to corroborate this (see Figure 4).
@@ 3.1. The urn connection
Let $p\in[0,1)$. For any $m\geq 1$, there exists a constant $C^{\prime}_{p,d,m}>0$ such that for any $a\in\mathcal{S}$, one has
$\mathbb{E}\bigg|\frac{N_{n}(a)}{n}-\frac{1}{d}\bigg|^{m}\leq(C^{\prime}_{p,d,m})^{m}\cdot r_{n}^{-m}.$
In fact, for any $m\geq 2$, one may take $C^{\prime}_{p,d,m}=C^{\prime}_{p,d}\cdot m$ for some constant $C^{\prime}_{p,d}>0$.
For the MERW on $\mathbb{Z}^{d}$, a result of this form was worked out in Lemma 4.3 of [Qin25a] for the moments of $\big|\frac{N_{n}(\mathbf{e}_{i})+N_{n}(-\mathbf{e}_{i})}{n}-\frac{1}{d}\big|$. The proof utilises the Burkholder-Davis-Gundy inequality [Dav70]. Since that argument works here almost verbatim, we omit the proof. We are explicitly specifying the dependence of $C^{\prime}_{p,d,m}$ on $m$ (see, e.g., [CT03, Chapter 11]) as we shall choose an optimal $m$ in a later argument.
@@
Let $p\in[0,1)$ and $m\geq 1$. There is a constant $\widetilde{C}_{p,d,m}>0$, such that
$\mathbb{E}|\Xi_{n}|^{m}\leq(\widetilde{C}_{p,d,m})^{m}\cdot r_{n}^{-m}.$
In fact, for any $m\geq 2$, one may take $\widetilde{C}_{p,d,m}=\widetilde{C}_{p,d}\cdot m$ for some constant $\widetilde{C}_{p,d}>0$. Consequently, for any $u\geq\frac{2\widetilde{C}_{p,d}\exp(1)}{r_{n}}$, one has
$\mathbb{P}(|\Xi_{n}|\geq u)\leq\exp\bigg(-\frac{ur_{n}}{\widetilde{C}_{p,d}\exp(1)}\bigg).$
@@
$\bigg|\frac{1}{n}\sum_{k=0}^{n-1}\bigg(\frac{\ell_{k}}{k}-\frac{\mathds{1}(\Delta_{k}>0)}{d}\bigg)\bigg|$ $\leq\frac{1}{n}\sum_{k=0}^{n-1}\bigg|\frac{\ell_{k}}{k}-\frac{\mathds{1}(\Delta_{k}>0)}{d}\bigg|$
$\leq\frac{1}{n}\sum_{k=1}^{n-1}\sum_{a\in\mathcal{S}}\bigg|\frac{N_{k}(a)}{k}-\frac{1}{d}\bigg|$
$=\sum_{a\in\mathcal{S}}\frac{1}{n}\sum_{k=1}^{n-1}\bigg|\frac{N_{k}(a)}{k}-\frac{1}{d}\bigg|.$
Applying Minkowski’s inequality and Lemma 3.1, we have
$\bigg[\mathbb{E}\bigg|\frac{1}{n}\sum_{k=0}^{n-1}\bigg(\frac{\ell_{k}}{k}-\frac{\mathds{1}(\Delta_{k}>0)}{d}\bigg)\bigg|^{m}\bigg]^{\frac{1}{m}}$ $=\sum_{a\in\mathcal{S}}\frac{1}{n}\sum_{k=1}^{n-1}\bigg[\mathbb{E}\bigg|\frac{N_{k}(a)}{k}-\frac{1}{d}\bigg|^{m}\bigg]^{\frac{1}{m}}$
$\leq C_{p,d,m}\cdot\sum_{a\in\mathcal{S}}\frac{1}{n}\sum_{k=1}^{n-1}r_{k}^{-1}.$
By applying Lemma 3.2 to the right-hand side, we conclude that for some constant $\widetilde{C}_{p,d,m}>0$, one has
$\big[\mathbb{E}|\Xi_{n}|^{m}\big]^{\frac{1}{m}}\leq\widetilde{C}_{p,d,m}\cdot r_{n}^{-1}.$
It is clear from the above proof that $\widetilde{C}_{p,d,m}=C^{\prime}_{p,d,m}\cdot C^{\prime\prime}_{p,d}$ for some constant $C^{\prime\prime}_{p,d}>0$. As such for any $m\geq 2$, we may take $\widetilde{C}_{p,d,m}=\widetilde{C}_{p,d}\cdot m$ for some constant $\widetilde{C}_{p,d}>0$.
Now, with $m\geq 2$,
$\mathbb{P}(|\Xi_{n}|\geq u)$ $\leq\frac{\mathbb{E}|\Xi_{n}|^{m}}{u^{m}}\leq\bigg(\frac{\widetilde{C}_{p,d}\cdot m}{ur_{n}}\bigg)^{m}\leq\sup_{t\geq 2}\bigg(\frac{\widetilde{C}_{p,d}\cdot t}{ur_{n}}\bigg)^{t}.$
It is easy to check that the supremum is attained at $t_{*}=\frac{ur_{n}}{\widetilde{C}_{p,d}\exp(1)}\geq 2$, which gives
$\mathbb{P}(|\Xi_{n}|\geq u)\leq\exp\bigg(-\frac{ur_{n}}{C_{p,d}\exp(1)}\bigg).$
This completes the proof. ∎